OpenStudy (anonymous):

!!!!!!Solve for x. 2x2-7x = 4

OpenStudy (anonymous):

2x^2-7x-4=0 (2x+1)(x-4)=0 2x+1=0 x=-1/2 or x-4=0 x=4

OpenStudy (anonymous):

these are the choices

OpenStudy (uri):

@14turnerk So much googling yo!

OpenStudy (anonymous):

so to solve it all x eguals 4 because if you plug it in the answer come back as 4

OpenStudy (anonymous):

Whitch option!!?!?!?!!!

OpenStudy (anonymous):

C